# The Value of Information in Human-AI Decision-making

### v1

@@ Abstract
Humans and AIs are often paired on decision tasks with the expectation of achieving complementary performance, where the combination of human and AI outperforms either one alone. However, how to improve performance of a human-AI team is often not clear without knowing more about what particular information and strategies each agent employs. We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflow. We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design. We propose a novel information-based instance-level explanation technique that adapts a conventional saliency-based explanation to explain information value in decision making.
@@ 1 Introduction
As the performance of artificial intelligence (AI) models improves, workflows in which human and AI model-based judgments are combined to make decisions are sought in medicine, finance, and other domains. Though statistical models often make more accurate predictions than human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954), whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024b).
@@
However, evidence of complementary performance between humans and AI is limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures like posthoc decision accuracy. This approach is problematic for several reasons. First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Second, it cannot provide insight into the potential for available information to improve the decisions, making it difficult to design interventions that improve the team’s performance.
@@
In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides more actionable information for intervening to improve the decision pipeline. For example, if human experts are found to possess decision-relevant information over the AI, we might collect further data to improve the AI model. If the model predictions contain decision-relevant information not contained in human decisions, we might design more targeted explanations to help humans integrate under-exploited information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible to in theory incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem. Based on the intuition that any information that is used by the agents will eventually reveal itself through variation in their decisions, we identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over the data-generating distribution. The second—instance-level human-complementary information value—supports analyses of how humans or AI systems use information on a instance-by-instance basis.
@@
We apply the framework by applying it to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is availabel at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018). First, we demonstrate its utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems. Next, we use our framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps human exploit available information for decision-making. Lastly, we demonstrate use of the framework to design explanations by extending SHAP Lundberg and Lee (2017) to highlight the portion of an AI’s prediction that complements human information.
@@ 2 Related work
Human-AI complementarity
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that on average, human–AI teams perform worse than the better of either humans or AI alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including whether the human and AI decisions provide additional information that is not used by the other.
Evaluation of human decision-making with machine learning
Our work contributes to the development of methods for evaluating decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) first proposed that evaluations of human-AI collaboration should be based on what information is available at the time of decisions. Our work contributes to definition of Bayesian best-attainable-performance benchmarks (Hofman et al., 2021; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) use a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on expected performance of any human-AI team. This benchmark provides a basis for identifying informational “opportunities” within a decision problem.
Human information in machine learning
One popular approach for developing machine learning models is to incorporate human information or expertise in model predictions (Alur et al., 2024a, b; Corvelo Benz and Rodriguez, 2023; Mozannar et al., 2024a; Bastani et al., 2021; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise but targets a broader scope by quantifying the information value for all available signals and agent decisions implied in a human–AI decision pipeline, rather than focusing solely on improving model performance.
@@ 3 Methodology
Our framework takes input as a decision problem associated with an information model and outputs the value of information in any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. Our framework provides two separate functions to quantify the value of information globally across the data-generating process and locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables to compare the agent-complementary information in signals for all possible decision problems.
Decision Problem
@@
Information Model
@@
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. These represent information obtained by a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temprature Celsius, etc. The decision-maker observes a signal, which is a combination of the basic signals, represented as a set $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day. Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{m}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$. The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$. Though $V$ is initially defined as a set of random variables, we will slightly abuse notation $V$ to represent a random variable that is drawn from the joint distribution of the basic signals in it.
@@
  • Data-generating process. A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. However, the DM may only observe a subset $V$ of the $n$ basic signals. We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
    $\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
@@
    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over unobserved signals.
Information value
@@
Our framework quantifies the value of information in a signal $V$ as the extent to which the payoff could be improved by the ideal use of $V$ over a baseline information set. We suppose a rational Bayesian DM who knows the data-generating process, observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff based on the posterior belief. Formally, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{v\sim\pi}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|v)}\![{S(d,\omega)}]}]$
@@
We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who has no access to a signal but only uses their prior belief to make decisions. In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$
@@
$\mathrm{R}^{\pi,S}(\emptyset)$ defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the information value of $V$ as
$IV^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$
We adopt the same idea to define the agent-complementary information values in our framework.
@@ 3.1 Agent-Complementary Information Value
Given the above definitions, it becomes possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their behaviors. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi^{b}\in\Delta(\mathbf{\Omega}\times\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi^{b}$ is given by the function:
$\mathrm{R}(D^{b})=\text{\bf E}_{d^{b}\sim\pi^{b}}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|D^{b}=d^{b})}\![{S(d,\omega)}]}]$
We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ and the information value in $D^{b}$, which we define as agent-complementary information value.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value of $V$ on agent decisions $D^{b}$ as
$ACIV^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$
If the $ACIV$ of a signal $V$ is low, this means either that the $IV$ of $V$ is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$). If, however, the $ACIV$ of $V$ is high, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
Furthermore, $ACIV$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the existing agent signal $D^{b}$, a large $ACIV$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans contribute on top of the AI. We demonstrate further usage of $ACIV$ in Section 4 and Section 5.
3.2 Instance-level Agent-Complementary
Instance-level Agent-Complementary Information Value ($ILIV$) evaluates the additional information contributed by a single realization of a signal rather than the entire joint distribution. This finer-grained view is critical for tasks where we need to understand the information value on individual instances, such as when asking whether a human expert should trust an individual prediction from an AI or how to help a human expert understand how the AI model is exploiting information about an instance.
To quantify the information value of a realization $V=v$, we construct $Z_{v}$ as a binary variable indicating whether signal $V$ is realized as $v$, i.e., $Z_{v}=\mathbb{1}[V=v]$. The data-generating model defining the joint distribution of $Z_{v}$, $\pi_{v}\in\Delta(\mathbf{\Omega}\times\{0,1\})$, can be constructed through transforming the original data-generating model $\pi$. Because $Z_{v}$ is a garbling of $V$ (Blackwell et al., 1951), there always exists a a Markov matrix $\Gamma:\mathbf{V}\rightarrow\{0,1\}$ such that the new data-generating process $\pi_{v}$ can be constructed through $\pi_{v}=\Gamma\pi$.

### v3

@@ Abstract
Multiple agents—including humans and AI models—are often paired on decision tasks with the expectation of achieving complementary performance, where the combined performance of both agents outperforms either one alone. However, knowing how to improve the performance of a human-AI team is often difficult without knowing more about what particular information and strategies each agent employs. We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows. We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design. We propose a novel information-based explanation technique that adapts SHAP, a saliency-based explanation to explain information value in decision making.
@@ 1 Introduction
As the performance of artificial intelligence (AI) for domain-specific predictions improves, workflows in which human and AI models are combined to make decisions in medicine, finance, and law, among other domains. Often, statistical models can be developed that exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954). However, whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024).
@@
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures based on post hoc decision accuracy (Passi and Vorvoreanu, 2022). This approach is problematic for several reasons. First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Second, it cannot provide insight into the potential for available information to improve decision performance, making it difficult to design interventions that improve the team’s performance.
@@
In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides actionable information for improving the decision pipeline. For example, the degree to which a model provides information that complements human judgments can guide AI model selection. Evidence that human experts possess decision-relevant information over AI predictions motivates further data collection to improve the AI model. Evidence that model predictions contain decision-relevant information that humans do not exploit can guide the design of explanations highlighting complementary information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem. The intuition behind this approach is that any information that is used by the agents will eventually reveal itself through variation in their decisions. We identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker. While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information.
@@
We demonstrate the framework by applying the metrics to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is available at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018). First, we demonstrate the framework’s utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems. Next, we use the framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps humans exploit available information for decision-making. Lastly, we demonstrate use of the framework to design explanations by extending SHAP (Lundberg and Lee, 2017) to highlight the portion of an AI’s explanation that complements human information.
@@ 2 Related work
Human-AI complementarity.
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
Evaluation of human decision-making with machine learning.
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of decisions. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
Human information in machine learning.
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a human–AI decision pipeline.
@@ 3 Methodology
Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
Decision Problem.
@@
In decision problems corresponding to prediction tasks, the decision space is a probabilistic belief over the state space, i.e., $\mathbf{D}=\Delta(\mathbf{\Omega})$. For such problems, a payoff function is said to be a proper scoring rule if the optimal action is to predict the true distribution, i.e., $p=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$. For any decision problem with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, there is an equivalent proper scoring rule $\hat{S}:\Delta(\mathbf{\Omega})\times\mathbf{\Omega}\to\mathbb{R}$ defined by choosing the optimal decision under the reported belief. Formally,
$\hat{S}(p,\omega)=S(\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}],\omega).$ (1)
Equation 1 shows a reduction from the payoff function $S$ to the proper scoring rule $\hat{S}$, i.e., any decision $p$ under $\hat{S}$ represents a decision $d$ in $S$ that best-responds to the distribution $p$. Therefore, the best-attainable performance defined in the proper scoring rule $\hat{S}$ is equivalent to the best-attainable performance defined in any payoff function $S$ that can be reduced to $\hat{S}$. Throughout our framework, we prefer proper scoring rules over non-proper scoring rules, since the best-attainable performance defined in the former implies the best-attainable performance in the latter.
Information Model.
@@
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temperature in Celsius, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. Specifically, we use $V=\{\Sigma_{j_{1}},\ldots,\Sigma_{j_{k}}\}$ for a signal having $k$ basic signals and denote the signal space as $\mathbf{V}=\mathbf{\Sigma}_{j_{1}}\times\ldots\times\mathbf{\Sigma}_{j_{k}}$. For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day. Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{k}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$. The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$.
@@
  • Data-generating process. A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. $\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals. The DM may only observe a subset $V$ of the $n$ basic signals. We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
    $\pi(\omega|v):=\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
@@
    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals.
The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process. When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process. In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure. These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images). Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings. We introduce an algorithm for identifying an “optimal” reduction model in decision problems in Appendix B. We demonstrate these two methods in Section 4 and Section 5 respectively.
Information value.
@@
Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set. We suppose a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief. Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (2)
where $d^{r}(\cdot):\mathbf{V}\rightarrow\mathbf{D}$ denotes the decision rule adopted by the rational DM.
$d^{r}(v)=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d,\omega)}]$ (3)
@@
We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions. In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (4)
@@
This baseline defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, the information value of $V$ is defined as
$\operatorname{IV}^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$ (5)
We adopt the same idea to define the agent-complementary information value in our framework.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
$\mathrm{R}^{\pi,S}(D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(d^{r}(d^{b}),\omega)}]$
We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ over the information value in $D^{b}$, which we define as the agent-complementary information value.
@@
Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value ($\operatorname{ACIV}$) of $V$ on agent decisions $D^{b}$ as
$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$ (6)
If the $\operatorname{ACIV}$ of a signal $V$ is small relative to the baseline (4), this means either that the information value of $V$ to the decision problem is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ or equivalent information to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$). If, however, the $\operatorname{ACIV}$ of $V$ is large relative to the baseline, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions. We demonstrate uses of $\operatorname{ACIV}$ in Section 4 and Section 5.
3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions. Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal. To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$. Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v$ on instances where $V=v$ is
$\mathrm{r}^{v,\pi,S}(v)=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v),\omega)}]$ (7)
where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v$ as defined in Equation 3. If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
$\mathrm{r}^{v,\pi,S}(v;D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v\cup d^{b}),\omega)}]$ (8)